OpenStudy (jjuden):

Is triangle ABC with vertices A(-1,4), B(3,1), and C(0,-3) a right triangle? Explain your answer using the slopes of the segments that form the sides of the triangle.

OpenStudy (princeharryyy):

slope ab = 1-4 / 3 +1 = -3/4 slope bc = -3-1 / 0 -3 = +4/3 slope ca = 4+3/-1 = -7 slope ab*bc = -1 which is the condition for right angles and using no other two slopes u get this condition fulfilled so triangle ab is a right angle triangle with right angle at b @jjuden